OpenStudy (anonymous):

how to write the equation of each line in slope -intercept form ...through (4,5) with slope -1/2

OpenStudy (anonymous):

y - y1 = m (x - x1) U have m = -1/2 and x1 = 4 and y1 = 5 plug in the values and u get the answer :)